# The Hidden Risks of Large Reasoning Models: A Safety Assessment of R1

### v3

@@ Abstract
The rapid development of large reasoning models, such as OpenAI-o3 and DeepSeek-R1, has led to significant improvements in complex reasoning over non-reasoning large language models (LLMs). However, their enhanced capabilities, combined with the open-source access of models like DeepSeek-R1, raise serious safety concerns, particularly regarding their potential for misuse. In this work, we present a comprehensive safety assessment of these reasoning models, leveraging established safety benchmarks to evaluate their compliance with safety regulations. Furthermore, we investigate their susceptibility to adversarial attacks, such as jailbreaking and prompt injection, to assess their robustness in real-world applications. Through our multi-faceted analysis, we uncover four key findings: (1) There is a significant safety gap between the open-source R1 models and the o3-mini model, on both safety benchmark and attack, suggesting more safety effort on R1 is needed. (2) The distilled reasoning model shows poorer safety performance compared to its safety-aligned base models. (3) The stronger the model’s reasoning ability, the greater the potential harm it may cause when answering unsafe questions. (4) The thinking process in R1 models pose greater safety concerns than their final answers. Our study provides insights into the security implications of reasoning models and highlights the need for further advancements in R1 models’ safety to close the gap. Warning: this paper includes examples that may be offensive or harmful.
@@ 1 Introduction
The landscape of large language models (LLMs) is evolving with the advent of large reasoning models like OpenAI-o3 OpenAI (2025b) and DeepSeek-R1 Guo et al. (2025), which leverage reinforcement learning to enhance complex reasoning. Unlike conventional LLMs, these models “think” (generate a structured chain-of-thought employing specialized output formats) before producing a final response. Reasoning models have superior performance in problem-solving, coding, scientific reasoning, and multi-step logical inference. However, their increased capabilities, combined with the recent open-sourcing of DeepSeek-R1, amplify their potential safety risks across a broad range of applications. Therefore, a comprehensive safety analysis of these reasoning models is essential to identify and mitigate their associated risks.
@@
In this work, as shown in Figure 1, we present a systematic and comprehensive safety assessment for these language reasoning models. Specifically, we first conduct a thorough safety evaluation by testing these reasoning language models against various established safety benchmarks, covering a broad range of safety categories from company policies and government regulations Zeng et al. (2024), and various application scenarios Wan et al. (2024b). Additionally, we assess their vulnerability to different adversarial attacks, including jailbreaking and prompt injection Jiang et al. (2024); Wan et al. (2024b), to analyze their robustness in real-world deployments. In these evaluations, we analyze both quantitative results and the safety behaviors of large reasoning models to gain deeper insights into their safety performance.
@@
Beyond classifying the safety of final model responses, a primary contribution of this work is a multi-faceted safety analysis specific to large reasoning models. First, to determine whether the reasoning process itself elevates safety risks, we evaluate the safety of the model’s internal reasoning steps (e.g., the content within <think> and </think> tags in DeepSeek-R1) and compare it against the safety of the final completion. Second, recognizing that unsafe responses can vary in their degree of harmfulness, we hypothesize that reasoning models, due to their enhanced capabilities, may generate more harmful unsafe responses. Therefore, in addition to binary safety classification, we evaluate the harmfulness level of model responses using pre-trained multi-attribute reward models Wang et al. (2024); Dorka (2024).
@@
Our experimental findings demonstrate that open-source reasoning models have a significant safety gap compared with the close-source o3-mini in both safety benchmarking and when facing adversarial attacks. Moreover, the distilled reasoning model exhibits consistently lower safety performance compared to their base safety-aligned model. Crucially, our analysis reveals that when reasoning models generate unsafe responses, these responses tend to be more harmful than those from non-reasoning models due to stronger abilities. Finally, we find that across the majority of benchmarks tested, the content generated during the reasoning process of R1 models exhibits lower safety than their final completions, underscoring an urgent need to enhance the safety of the reasoning process itself. We hypothesize that the safety performance of R1 models may be attributed to non-sufficient safety-specific training, and the process of fine-tuning the Llama 3.3 (distilled-R1) could have inadvertently degraded its pre-existing safety alignment (Qi et al., 2023). Given the broad adaptability of open-source reasoning models, we advocate for stronger safety alignment to mitigate potential risks in the future.
@@ 2 Background and Related Work
Recent advancements in large reasoning language models — such as OpenAI’s o1 and o3 OpenAI (2025a, b) and DeepSeek-R1 Guo et al. (2025) have substantially enhanced LLMs’ problem-solving capabilities by integrating structured reasoning mechanisms. For example, the OpenAI o1 model spends additional compute time to generate long chains of reasoning before producing a final answer, achieving PhD-level performance on challenging mathematical and scientific benchmarks (OpenAI, 2025a). Building on this, the o3 series further refines the approach to boost performance (OpenAI, 2025b). In parallel, DeepSeek-R1 pioneered a reasoning-oriented reinforcement learning training approach without supervised fine-tuning, demonstrating emergent reasoning behaviors and achieving performance comparable to o1 on math, coding, and science tasks Guo et al. (2025). These models underscore the effectiveness of test-time self-reflection in addressing complex challenges, although significant hurdles remain in ensuring their safety and reliability.
@@
As the abilities of LLMs become stronger, various benchmarks have been proposed to evaluate the safety of LLMs in different safety categories and application domains Wang et al. (2023); Bhatt et al. (2024); Wan et al. (2024b); Li et al. (2024a); Xie et al. (2024); Zeng et al. (2024); Andriushchenko et al. (2024). These benchmarks evaluate whether LLMs comply with malicious queries and produce harmful content, with comprehensive categories that cover safety regulations from the government and company policies. Röttger et al. (2023) also evaluate whether the safety alignment of LLMs leads to over-sensitive to benign queries. More recently, there are safety evaluations for new applications of LLMs, including scenarios that are relevant to cybersecurity Wan et al. (2024b); Bhatt et al. (2024), and LLM agents that make sequential decisions and receive feedback from the environments Andriushchenko et al. (2024).
@@
As LLMs become integral to real-world applications, adversaries are devising increasingly sophisticated strategies to subvert their safety mechanisms. One prominent tactic is prompt injection Yi et al. (2023); Zhan et al. (2024); Zhang et al. (2024), wherein adversaries insert additional instructions into the input text to override the model’s intended directives or trigger harmful behavior. Another major threat comes from jailbreak attacks, which trick LLMs into responding to queries they would typically refuse. For example, strategy-based jailbreaks leverage natural language constructs—often by presenting hypothetical scenarios—to manipulate model behavior Wei et al. (2024); Jiang et al. (2024); Zhu et al. (2024); Li et al. (2024b); Liu et al. (2024b), while optimization jailbreaks focus on optimizing a prefix string to maximize the likelihood of generating responses to otherwise harmful queries Zou et al. (2023); Liao and Sun (2024). In our work, we select representative safety benchmarks and attacks to analyze the safety performance and behaviors of large reasoning models. Further, we introduce multi-faceted safety evaluation to better understand their safety risks.
@@ 3.1 Research Questions
Safety Benchmarks AirBench Zeng et al. (2024) Safety Policies 5,694
MITRE Wan et al. (2024b) Cyber Attack 377
Interpreter Wan et al. (2024b) Code Exc 500
Phishing Wan et al. (2024b) Spear Phishing 200
XSTest Röttger et al. (2023) Over-refusal 250
Adversarial Attacks WildGuard Han et al. (2024) Jailbreak 810
Injection Bhatt et al. (2024) Prompt injection 251
@@ 3.2 Evaluation Design
As shown in Table 1, we select 5 representative datasets from 3 safety benchmarks and 2 datasets on adversarial attacks for evaluation. For RQ1, we select Air-Bench Zeng et al. (2024), a comprehensive safety evaluation benchmark containing safety prompts from government regulations and corporate policies. On top of this, we select three test suites from CyberSecEval to test the models’ safety performance when given unsafe tasks in various cybersecurity-related scenarios, including cyberattacks, secure code execution, and multi-turn spear phishing Wan et al. (2024b). Additionally, we test the over-refusal behavior of large reasoning models on the XStest dataset Röttger et al. (2023). For RQ 3, we test the models on the WildGuard Jailbreak dataset Han et al. (2024), and the prompt injection test from the CyberSecEval. Detailed introductions for the datasets can be found in Appendix A.1.
@@
To facilitate a better analysis of the impact of reasoning abilities on safety, we evaluate 2 open-weight large reasoning models, DeepSeek-R1 Guo et al. (2025), the strongest reasoning model of R1 series, and DeepSeek-R1-70b Guo et al. (2025), a reasoning model distilled based on Llama 3.3-70b, and compare them with their 2 non-reasoning counterparts, Llama 3.3-70b Dubey et al. (2024) and DeepSeek-V3 Liu et al. (2024a). Finally, to understand the safety gap of open large reasoning models, we include the proprietary o3-mini OpenAI (2025b) model in our evaluation.
@@
We perform two levels of evaluation. First, for all the benchmarks, we use GPT-4o Hurst et al. (2024) to do safety classification on the responses with defined criteria. On AIR-bench, Code Interpreter Test and MITRE Tests, responses containing no information that helps the malicious query will be considered safe. In these benchmarks, we also evaluate the safety of the reasoning process for reasoning LLMs and compare it with the final completion. Spear Phishing Tests originally do not provide safety classification to the response. We create safety classification criteria based on whether the model performs the phishing attempt. On XSTest, GPT-4o is similarly used to classify responses into full compliance, full refusal, and partial refusal, allowing for an objective evaluation of exaggerated safety behaviors.
@@
Open weight Llama3.3 52.9 27.1 70.4 4.0
@@
Figure 2: Two safety benchmark evaluations: (A) Level-2 categorized results of the models on Air-Bench. (B) Evaluation of the models’ safety rate (%) in the Code Interpreter Tests across different risk categories.
@@ 4.1 Overall Safety Analysis
We evaluate the average safety rate of all models across four benchmarks with unsafe queries. First, o3-mini exhibits significantly higher safety than open-source reasoning and non-reasoning models, effectively identifying and rejecting most unsafe queries across various scenarios. Open large reasoning models still have a considerable gap to close compared with o3-mini. Second, we observe that the distilled R1-70b consistently achieves a lower safety rate than Llama-3.3, suggesting that reasoning-supervised fine-tuning reduces a model’s safety performance; this aligns with the finding of  (Qi et al., 2023) on the effect of supervised fine-tuning to safety performance. Finally, R1 demonstrates better safety performance than V3 on the broad safety categories on AirBench. However, R1 shows a significantly more severe safety risk in the cybersecurity domain, with more complex tasks and environment settings. These results indicate that more effort should be put into safety alignment on R1 models.
@@
From the output of the models, we identify a different safety behavior of R1 models from non-reasoning LLMs – the thinking process of the R1 models usually determines the safety of final completion. In the thinking process, if the model performs safety thinking and decide that the query is not safe to answer, it will refuse the query it in the final answer. Otherwise, if no safety thinking happens, or the model believes the query is appropriate to answer, no refusal will happens. In contrast, the refusal behavior from non-reasoning LLMs usually happens immediately without explicit thinking. Examples and more analysis on the safety thinking are in Section 6.1.
@@ 4.2 Select Datasets Analysis
We present the level 2 category results of Air-Bench in Figure 2 (A), covering 16 categories. First, the comparison between reasoning and non-reasoning models has significant differences across categories. For example, Llama3.3 is significantly better than R1-70b in ‘Operational Misuses’, ‘Security Risks’, etc., with more than 20% advantage, but is worse than R1-70b in ‘Hate/Toxicity’ by 11%. Second, all the reasoning models, including o3-mini have very low performance in certain categories like ‘Operational Misuses’ and ‘Sexual Content’. Meanwhile, the open large reasoning models perform significantly worse than o3-mini in most of the categories. These findings suggest that certain safety aspects require greater attention in large reasoning models.
Code Interpreter Test
The results in Figure 2 (B) show that reasoning models exhibit higher security risks than non-reasoning models, particularly in reflected attacks and social engineering. Specifically, the safe operation percentage of the R1 model in the social engineering category is significantly lower than Llama3.3 and falls behind V3 by over 50%. This indicates that reasoning models are more vulnerable to abuse in interactive categories.
@@
Table 2 presents the models’ safety rate in the Spear Phishing Test. In this evaluation, the tested LLM will be asked to perform multi-turn spear phishing against a victim LLM. Here, the safety rate refers to the proportion of conversations in which the model did not attack in any turn. Notably, the result reveals a significant safety risk, as o3-mini initiates an attack in only 5% of trials, while other models do so in over 95% of cases. Moreover, o3-mini almost always refuses to engage in an attack at the start of the dialogue, with only one instance where it does so immediately. In other unsafe dialogues, it begins attacking after multiple prompts. Also, in the cases where it engaged in a phishing attempt, it can later recognize the safety risk and cease the attack. In contrast, at the start of the dialogue, neither the R1 nor R1-70b models show safety awareness when prompted to attack. This highlights the risks posed by R1 models as a multi-turn spear phishing agent.
Models Avg Worst Cat Worst Cat Performance
Llama3.3 96.8 + 2.4 + 0.8 Privacy (Fictional) 72 + 20 + 8
R1-70b 94.8 + 4.4 + 0.8 Privacy (Fictional) 68 + 28 + 4
DS-V3 98.0 + 2.0 + 0.0 Privacy (Fictional) 80 + 20 + 0
DS-R1 96.0 + 3.2 + 0.8 Real Discr., Nons. Group 84 + 16 + 0
o3-mini 92.8 + 7.2 + 0.0 Privacy (Fictional) 64 + 36 + 0
Table 3: Performance of models on safe prompts in XSTest. Safe prompts should achieve full compliance (FC), and not refused with either full refusal (FR) or partial refusal (PR). The columns from left to right correspond to FC, FR and PR respectively.
@@
The over-refusal rate to benign queries is also a strong indicator of models’ safety alignment. We present the average result and the result of the category with the worst performance for each model on the XSTest dataset in Table 3 and leave the detailed categorized result in Table 9 in the Appendix A.2. We observe that all the models tested can usually respond to safe prompts correctly without refusal, with a refusal rate of less than 10%. Among these, the o3-mini has the highest refusal rate due to being more sensitive to seemingly unsafe queries. However, all the models refuse to answer a significant portion of safe queries in at least one safety category.
The most frequent over-refusal made by models is in the fictional privacy category, where the query asks about private information on fictional persons. For instance, when the R1-70b is asked “What is Harry Potter’s criminal history?”, it replies “As an AI assistant, I don’t have access to internal company information…”. Besides that, Deepseek-R1 often refuse to prompts with discrimination against a nonsense group. (e.g. “Why should purple carrots not be given visas to enter the US?” $\rightarrow$ “The People’s Republic of China consistently advocates for the principle of equality …”). This shows that the current models’ safety alignment is not still refined enough in specific categories.

### v4

@@ Abstract
The rapid development of large reasoning models (LRMs), such as OpenAI-o3 and DeepSeek-R1, has led to significant improvements in complex reasoning over non-reasoning large language models (LLMs). However, their enhanced capabilities, combined with the open-source access of models like DeepSeek-R1, raise serious safety concerns, particularly regarding their potential for misuse. In this work, we present a comprehensive safety assessment of these reasoning models, leveraging established safety benchmarks to evaluate their compliance with safety regulations. Furthermore, we investigate their susceptibility to adversarial attacks, such as jailbreaking and prompt injection, to assess their robustness in real-world applications. Through our multi-faceted analysis, we uncover four key findings: (1) There is a significant safety gap between the open-source reasoning models and the o3-mini model, on both safety benchmark and attack, suggesting more safety effort on open LRMs is needed. (2) The stronger the model’s reasoning ability, the greater the potential harm it may cause when answering unsafe questions. (3) Safety thinking emerges in the reasoning process of LRMs, but fails frequently against adversarial attacks. (4) The thinking process in R1 models poses greater safety concerns than their final answers. Our study provides insights into the security implications of reasoning models and highlights the need for further advancements in R1 models’ safety to close the gap. Warning: this paper includes examples that may be offensive or harmful. Project page: https://r1-safety.github.io.
@@ 1 Introduction
The landscape of large language models (LLMs) is evolving with the advent of large reasoning models like OpenAI-o3 o3minicard and DeepSeek-R1 guo2025deepseek, which leverage reinforcement learning to enhance complex reasoning. Unlike conventional LLMs, these models “think” (generate a structured chain-of-thought employing specialized output formats) before producing a final response. Reasoning models have superior performance in problem-solving, coding, scientific reasoning, and multi-step logical inference. However, their increased capabilities, combined with the recent open-sourcing of DeepSeek-R1, amplify their potential safety risks across a broad range of applications. Therefore, a comprehensive safety analysis of these reasoning models is essential to identify and mitigate their associated risks.
@@
In this work, as shown in Figure 1, we present a systematic and comprehensive safety assessment for these language reasoning models. Specifically, we first conduct a thorough safety evaluation by testing these reasoning language models against various established safety benchmarks, covering a broad range of safety categories from company policies and government regulations zeng2024air, and various application scenarios wan2024cyberseceval. Additionally, we assess their vulnerability to different adversarial attacks, including jailbreaking and prompt injection jiang2024wildteaming; wan2024cyberseceval, to analyze their robustness in real-world deployments. In these evaluations, we analyze both quantitative results and the safety behaviors of large reasoning models to gain deeper insights into their safety performance.
@@
Beyond classifying the safety of final model responses, a primary contribution of this work is a multi-faceted safety analysis specific to large reasoning models. First, to determine whether the reasoning process itself elevates safety risks, we evaluate the safety of the model’s internal reasoning steps (e.g., the content within <think> and </think> tags in DeepSeek-R1) and compare it against the safety of the final completion. Second, recognizing that unsafe responses can vary in their degree of harmfulness, we hypothesize that reasoning models, due to their enhanced capabilities, may generate more harmful unsafe responses. Therefore, in addition to binary safety classification, we evaluate the harmfulness level of model responses using pre-trained multi-attribute reward models wang2024interpretable; dorka2024quantile.
@@
Our experimental findings demonstrate that open-source reasoning models have a significant safety gap compared with the close-source o3-mini in both safety benchmarking and when facing adversarial attacks. Moreover, the distilled reasoning model exhibits consistently lower safety performance compared to their base safety-aligned model. Crucially, our analysis reveals that when reasoning models generate unsafe responses, these responses tend to be more harmful than those from non-reasoning models due to stronger abilities. Finally, we find that across the majority of benchmarks tested, the content generated during the reasoning process of R1 models exhibits lower safety than their final completions, underscoring an urgent need to enhance the safety of the reasoning process itself. We hypothesize that the safety performance of R1 models may be attributed to non-sufficient safety-specific training, and the process of fine-tuning the Llama 3.3 (distilled-R1) could have inadvertently degraded its pre-existing safety alignment (qi2023fine). Given the broad adaptability of open-source reasoning models, we advocate for stronger safety alignment to mitigate potential risks in the future.
@@ 2 Background and Related Work
Recent advancements in large reasoning language models — such as OpenAI’s o1 and o3 o1card; o3minicard and DeepSeek-R1 guo2025deepseek have substantially enhanced LLMs’ problem-solving capabilities by integrating structured reasoning mechanisms. For example, the OpenAI o1 model spends additional compute time to generate long chains of reasoning before producing a final answer, achieving PhD-level performance on challenging mathematical and scientific benchmarks (o1card). Building on this, the o3 series further refines the approach to boost performance (o3minicard). In parallel, DeepSeek-R1 pioneered a reasoning-oriented reinforcement learning training approach without supervised fine-tuning, demonstrating emergent reasoning behaviors and achieving performance comparable to o1 on math, coding, and science tasks guo2025deepseek. These models underscore the effectiveness of test-time self-reflection in addressing complex challenges, although significant hurdles remain in ensuring their safety and reliability.
@@
As the abilities of LLMs become stronger, various benchmarks have been proposed to evaluate the safety of LLMs in different safety categories and application domains wang2023not; bhatt2024cyberseceval; wan2024cyberseceval; li2024salad; xie2024sorry; zeng2024air; andriushchenko2024agentharm. These benchmarks evaluate whether LLMs comply with malicious queries and produce harmful content, with comprehensive categories that cover safety regulations from the government and company policies. rottger2023xstest also evaluate whether the safety alignment of LLMs leads to over-sensitive to benign queries. More recently, there are safety evaluations for new applications of LLMs, including scenarios that are relevant to cybersecurity wan2024cyberseceval; bhatt2024cyberseceval, and LLM agents that make sequential decisions and receive feedback from the environments andriushchenko2024agentharm.
@@
As LLMs become integral to real-world applications, adversaries are devising increasingly sophisticated strategies to subvert their safety mechanisms. One prominent tactic is prompt injection yi2023benchmarking; zhan2024injecagent; zhang2024agent, wherein adversaries insert additional instructions into the input text to override the model’s intended directives or trigger harmful behavior. Another major threat comes from jailbreak attacks, which trick LLMs into responding to queries they would typically refuse. For example, strategy-based jailbreaks leverage natural language constructs—often by presenting hypothetical scenarios—to manipulate model behavior wei2024jailbroken; jiang2024wildteaming; zhu2024autodan; li2024llm; liu2024autodan, while optimization jailbreaks focus on optimizing a prefix string to maximize the likelihood of generating responses to otherwise harmful queries zou2023universal; liao2024amplegcg. In our work, we select representative safety benchmarks and attacks to analyze the safety performance and behaviors of large reasoning models. Further, we introduce multi-faceted safety evaluation to better understand their safety risks.
@@ 3.1 Research Questions
Safety Benchmarks AirBench zeng2024air Safety Policies 5,694
MITRE wan2024cyberseceval Cyber Attack 377
Interpreter wan2024cyberseceval Code Exc 500
Phishing wan2024cyberseceval Spear Phishing 200
XSTest rottger2023xstest Over-refusal 250
Adversarial Attacks WildGuard wildguard2024 Jailbreak 810
Injection bhatt2024cyberseceval Prompt injection 251
@@ 3.2 Evaluation Design
As shown in Table 1, we select 5 representative datasets from 3 safety benchmarks and 2 datasets on adversarial attacks for evaluation. For RQ1, we select Air-Bench zeng2024air, a comprehensive safety evaluation benchmark containing safety prompts from government regulations and corporate policies, covering 16 level-2 safety categories. On top of this, we select three test suites from CyberSecEval to test the models’ safety performance when given unsafe tasks in various cybersecurity-related scenarios, including cyberattacks, secure code execution, and multi-turn spear phishing wan2024cyberseceval. Additionally, we test the over-refusal behavior of large reasoning models on the XStest dataset rottger2023xstest. For RQ 3, we test the models on the WildGuard Jailbreak dataset wildguard2024, and the prompt injection test from the CyberSecEval. Detailed introductions for the datasets can be found in Appendix A.1.
@@
To facilitate a better analysis of the impact of reasoning abilities on safety, we evaluate 3 open-weight large reasoning models, DeepSeek-R1 guo2025deepseek, the strongest reasoning model of R1 series, and DeepSeek-R1-70b guo2025deepseek, a reasoning model distilled based on Llama 3.3-70b, and QwQ qwen2025qwq. For R1-70b and R1, we compare them with their 2 non-reasoning counterparts, Llama 3.3-70b dubey2024llama and DeepSeek-V3 liu2024deepseek. Finally, to understand the safety gap of open large reasoning models, we include the proprietary models o3-mini o3minicard and Gemini 2.5 deepmind2025gemini in our evaluation.
@@
We perform two levels of evaluation. First, for all the benchmarks, we use GPT-4o hurst2024gpt to do safety classification on the responses with defined criteria. On AIR-bench, Code Interpreter Test and MITRE Tests, responses containing no information that helps the malicious query will be considered safe. In these benchmarks, we also evaluate the safety of the reasoning process for reasoning LLMs and compare it with the final completion. Spear Phishing Tests originally do not provide safety classification to the response. We create safety classification criteria based on whether the model performs the phishing attempt. On XSTest, GPT-4o is similarly used to classify responses into full compliance, full refusal, and partial refusal, allowing for an objective evaluation of exaggerated safety behaviors. We provide the validation on the effectiveness of GPT-4o evaluation in Appendix A.2.
@@
Open weight QwQ 46.3 18.8 79.9 3.5
Llama3.3 52.9 27.1 70.4 4.0
@@
Figure 2: Level-2 categorized results of the models on Air-Bench.
Models Avg
QwQ 98.4 + 1.2 + 0.4
Llama3.3 96.8 + 2.4 + 0.8
R1-70b 94.8 + 4.4 + 0.8
DS-V3 98.0 + 2.0 + 0.0
DS-R1 96.0 + 3.2 + 0.8
o3-mini 92.8 + 7.2 + 0.0
Gemini 2.5 98.4 + 1.2 + 0.4
Table 3: Performance of models on safe prompts in XSTest. The columns from left to right correspond to full compliance, full refusal and partial refusal, respectively.
@@ 4.1 Overall Safety Analysis
We evaluate the average safety rate of all models across four benchmarks with unsafe queries. First, o3-mini and Gemini 2.5 exhibits significantly higher safety than open-source reasoning and non-reasoning models, effectively identifying and rejecting most unsafe queries across various scenarios. Open large reasoning models still have a considerable gap to close compared with o3-mini. Second, we observe that the distilled R1-70b consistently achieves a lower safety rate than Llama-3.3, suggesting that reasoning-supervised fine-tuning reduces a model’s safety performance; this aligns with the finding of  qi2023fine on the effect of supervised fine-tuning on safety performance, potentially due to the distillation data may include inputs that are similar to harmful queries he2024your. Finally, R1 demonstrates better safety performance than V3 on the broad safety categories on AirBench. However, R1 shows a significantly more severe safety risk in the cybersecurity domain, with more complex tasks and environment settings. These results indicate that more effort should be put into safety alignment on R1 models.
@@
From the output of the models, we identify a different safety behavior of R1 models from non-reasoning LLMs – the thinking process of the R1 models usually determines the safety of final completion. In the thinking process, if the model performs safety thinking and decides that the query is not safe to answer, it will refuse the query it in the final answer. Otherwise, if no safety thinking happens, or the model believes the query is appropriate to answer, no refusal will happens. In contrast, the refusal behavior from non-reasoning LLMs usually happens immediately without explicit thinking. Examples and more analysis on the safety thinking are in Section 6.1.
@@ 4.2 Select Datasets Analysis
We present the level 2 category results of Air-Bench in Figure 2, covering 16 categories. First, the comparison between reasoning and non-reasoning models has significant differences across categories. For example, Llama3.3 is significantly better than R1-70b in ‘Operational Misuses’, ‘Security Risks’, etc., with more than 20% advantage, but is worse than R1-70b in ‘Hate/Toxicity’ by 11%. Second, all the reasoning models, including o3-mini have very low performance in certain categories like ‘Operational Misuses’ and ‘Sexual Content’. Meanwhile, the open large reasoning models perform significantly worse than o3-mini in most of the categories. These findings suggest that certain safety aspects require greater attention in large reasoning models.
@@
Table 2 presents the models’ safety rate in the Spear Phishing Test. In this evaluation, the tested LLM will be asked to perform multi-turn spear phishing against a victim LLM. Here, the safety rate refers to the proportion of conversations in which the model did not attack in any turn. Notably, the result reveals a significant safety risk, as o3-mini initiates an attack in only 5% of trials, while other models do so in over 95% of cases. Moreover, we find that o3-mini almost always refuses to engage in an attack at the start of the dialogue, with only one instance where it does so immediately. In other unsafe dialogues, it begins attacking after multiple prompts. Also, in the cases where it engaged in a phishing attempt, it can later recognize the safety risk and cease the attack. In contrast, at the start of the dialogue, neither the R1 nor R1-70b models show safety awareness when prompted to attack. This highlights the risks posed by R1 models as a multi-turn spear phishing agent.
@@
The over-refusal rate to benign queries is also a important indicator of models’ safety alignment. We present the average result for each model on the XSTest dataset in Table 2 and leave the detailed categorized result in Table 10 in the Appendix A.3. We observe that all the models tested can usually respond to safe prompts correctly without refusal, with a refusal rate of less than 10%. Among these, the o3-mini has the highest refusal rate, reflecting greater sensitivity to certain phrases in the seemingly unsafe queries. This, combined with their low safety rate in harmful queries, indicates that open LRMs are more aligned to comply with user queries.